Lenow’s Drug Stores and Hall’s Pharmaceuticals are competitors in the discount drug chain store business. The separate capital structures for Lenow and Hall are presented here.

Lenow Hall
Debt @ 8% $ 270,000 Debt @ 8% $ 540,000
Common stock, $10 par 540,000 Common stock, $10 par 270,000
Total $ 810,000 Total $ 810,000
Common shares 54,000 Common shares 27,000

a. Complete the following table given earnings before interest and taxes of $31,000, $64,800, and $81,000. Assume the tax rate is 10 percent. (Negative amounts should be indicated by parentheses or a minus sign. Round your answers to 2 decimal places.)
  

EBIT Total Assets EBIT/TA % Lenow EPS Hall EPS

31,000 $810,000

64,800 $810,000

81,000 $810,000
b-1. What is the EBIT/TA rate when the firm's have equal EPS?


b-2. What is the cost of debt?
  


b-3. State the relationship between earnings per share and the level of EBIT.
  


c. If the cost of debt went up to 10 percent and all other factors remained equal, what would be the break-even level for EBIT?
Answer #1

Please provide rating..

Answer a.
i ii iii=i/ii
EBIT Total asset EBIT/TA Lenow EPS Hall EPS
31000 810000 3.83%                          0.16 -0.41
64800 810000 8.00% 0.72 0.72
81000 810000 10.00% 0.99 1.26
Computation of EPS
Lenow Hall
i EBIT 31000 64800 81000 31000 64800 81000
ii interest 21600 21600 21600 43200 43200 43200
iii=i-ii EBT 9400 43200 59400 -12200 21600 37800
iv=iii*.7 PAT=EBT*(1-0.1) 8460 38880 53460 -10980 19440 34020
v Share out standing 54000 54000 54000 27000 27000 27000
vi=iv/v EPS              0.16          0.72                          0.99        (0.41)          0.72          1.26
Answer b.-1
at 8% EBIT/TA both the firm has equal EPS
Answer b.-2
Cost of debt =8%
Answer b.-3
relationship
Pretax return is lower than cost of debt therefore Hall EPS is lower compared to Lenow
Pretax return is equal to the cost of debt therefore both EPS are equal
pretax return is higher than cost of debt therefore Lenow EPS is lower than Hall EPS
Answer c
Break even level if debt cost incresed to 10%
At break even (with 10% cost of debt) we should have EBIT/TA = 10%
EBIT = 12%* TA
EBIT = 10%*810000
EBIT = 81000
Breakeven EBIT = 81000